OpenStudy (anonymous):

Help, Fan and medal. Assuming x ≠ 0 and y ≠ 0, what is the quotient of 8x^9y^3 + 2x^6y^2 + 8x^3y/2x^3y a- 4x^27y^3 + x^18y^2 + 4x^9y b - 4x^3y^3 + x^2y^2 + 4xy c - 4x^6y^2 + x^3y + 4 d - 4x^12y^4 + x^9y^3 + 4x^6y^2

OpenStudy (freckles):

\[\frac{8x^9y^3+2x^6y^2+8x^3y}{2x^3y} \\ =\frac{8x^9y^3}{2x^3y}+\frac{2x^6y^2}{2x^3y}+\frac{8x^3y}{2x^3y}\]

OpenStudy (freckles):

try to simplify each fraction

OpenStudy (anonymous):

I have no idea how :(